OpenStudy (anonymous):

f(x) = e^6x + e^−x Find increasing and decreasing intervals. I understand how to do these questions, but I get really confused when they do it with the e

OpenStudy (gorv):

ok how u find increasing interval??

OpenStudy (anonymous):

I need to find the derivative and then do a number line

OpenStudy (gorv):

u know the derivative of e^ something??

OpenStudy (anonymous):

So far I got up to =6e^6x-e^-e

OpenStudy (anonymous):

I dont know what to do after that

OpenStudy (gorv):

what u do to find increasing or decreasing interval in general ??

OpenStudy (gorv):

after finding derivative

OpenStudy (anonymous):

I found the critical numbers and then made a number line and saw where it was increasing and decreasing

OpenStudy (anonymous):

I set the first derivative = 0

OpenStudy (gorv):

\[6e^{6x}-e^{-x}=0\]

hartnn (hartnn):

The function is increasing when its first derivative is ---- The function is decreasing when its first derivative is ---- know what comes in those blanks ?

OpenStudy (gorv):

add e^-x on both side

OpenStudy (anonymous):

would I then get e^-x (6e^5x-1)=0 ?

OpenStudy (gorv):

hartnn said right first we need to assure it is greater than 0 or less than 0 after that we will find the interval by puttin it equal to zero

OpenStudy (gorv):

\[6e^{6x}-e^{-x}+e^{-x}=e^{-x}\]

OpenStudy (gorv):

what u will get ??

OpenStudy (anonymous):

6e^6x=e^-x

OpenStudy (gorv):

yeah ...u know?? x^(-a)=1/x^a it can also written like this also

OpenStudy (gorv):

@sheetalvee

OpenStudy (anonymous):

oh okay, I didnt know that

OpenStudy (gorv):

okay leave ...it ..take natural log on both side

OpenStudy (anonymous):

okay. so 6x log 6e = -x log 6 ?

OpenStudy (gorv):

\[\log_e (6*e^6x)=\log_e e^{-x}\]

OpenStudy (gorv):

log 6+ 6x*log e= -x log e

OpenStudy (gorv):

when number and base are equal the its value =1 like here base= e and number =e

OpenStudy (gorv):

log6+6x*1=-x*1 tell me what u got

OpenStudy (anonymous):

log6=-7x ?

OpenStudy (gorv):

log_e 6=1.79

OpenStudy (gorv):

1.79=-7x

OpenStudy (gorv):

divide both side by 7

hartnn (hartnn):

i prefer to take inequalities from beginning only increasing when \(6 e^{6x} > e^{-x} \) decreasing when \(6 e^{6x} < e^{-x} \)

OpenStudy (gorv):

yes thatz is the best way..i carried bcoz she said she put =0

OpenStudy (anonymous):

wait how did you get 1.79

OpenStudy (gorv):

u got the crictical point now

OpenStudy (gorv):

we took natural log so natural log of 6 = 1.79

hartnn (hartnn):

also by now you should be able to solve it like this :  without using logs if you can't, then better learn it now :)